OpenStudy (pphalke):

Find the total monthly payment for a $185,000 mortgage loan at 4.5% for 30 years. The assessed value of the home is $200,000. The annual taxes on the home are 0.9% of the assessed value, and the insurance on the home costs $375 per year I checked out the last problem and the answer was 1668.88 this one is the same but the numbers are different can I get help please

OpenStudy (mathmate):

Did you know how to get $1668.88 in the last problem? If you did, it's the same way here. This house will cost less per month, because less borrowing, less interest, mortgaged over 30 years instead of 20, less taxes, less insurance! It's only about three quarters of 1668.88!

OpenStudy (pphalke):

my book for this one says $1,118.62 but hw do I get to this answer what do I have to do? I know its the same way but I need steps

OpenStudy (pphalke):

@mathmate can I have some help on how to do the steps I know the formula

OpenStudy (mathmate):

Sure! You would use your formula to find the monthly payment for the mortgage of 185000. Tell me what you get!

OpenStudy (mathmate):

Then add on (taxes and insurance)/12 to get the monthly payments for these two. Practical note: Most company will charge interest when annual payments are paid by the month, but we'll ignore that.

OpenStudy (pphalke):

how do I plug the numbers in? I don't want to be wrong

OpenStudy (mathmate):

Do you use a calculator or a computer calculator like Google?

OpenStudy (mathmate):

do you use this formula from your last question? 

OpenStudy (pphalke):

yea and yea I did use that one plugging in and calculating I could be wrong

OpenStudy (mathmate):

So what did you get out of the formula?

OpenStudy (pphalke):

I haven't plugged in and I don't know where to put the numbers inside the formula

OpenStudy (mathmate):

I use a different formula, but this one works too. Here are the definitions of items for this formula: rate=interest per \(month\) in decimal, so for 4.5% per annum becomes 0.045/12=0.00375 principal=185000 (hope that's obvious to you) months = number of months = 30*12=360 (hope that's obvious to you too) The only catch is you have to be very careful with parentheses. ALWAYS put parentheses around both numerator and denominator to avoid unnecessary mistakes. There is no way to get around it but to practise!

OpenStudy (pphalke):

I got the first part what you said but if I ever did plug in I wouldn't ever have the answer of $1,118.62

OpenStudy (mathmate):

What did you get for the first part (mortgage). Just to check. The other two parts are much easier! lol